Using boolean algebra, simplify each expression:

_

(B + BC)(B + B^C)(B + D)


1
Expert's answer
2021-09-14T01:54:25-0400

AB + A'C + B'C

= AB + A'BC + A'B'C + B'C

= AB + A'BC + B'C (1 + A')

= AB + A'BC + B'C

= ABC + ABC' + A'BC + B'C

= BC(A+A') + ABC' + B'C

= BC + B'C + ABC'

= C + ABC'

I have a second question, might as well add it here because it's also simplification of boolean algebra.

f = cx + ac'x + bc'x + a'b'c'x' (used a K-map to generate this, now I have to simplify further)

f = c'x(a+b) + cx + a'b'c'x' - no idea how to continue from here